The chip sequence for each of the sources in a network are A[+1 +1 -1 -1], B[+1 -1 +1 -1], C[+1 +1 -1 -1], D[+1 -1 -1 +1]. The data recieved on the channel is [-1 -1 -3 +1]. Find out the data sent from B and C. Show your work.
Answer #1

Chip sequence of A = [+1 +1 -1 -1]

Chip sequence of B = [+1 -1 +1 -1]

Chip sequence of C = [+1 +1 -1 -1]

Chip sequence of D = [+1 -1 -1 +1]

Data on the channel =  [-1 -1 -3 +1]

To find the data sent by Source B, we take product (Scalar product / dot product) of  the data on the channel with the chip sequence of source B

[-1 -1 -3 +1] . [+1 -1 +1 -1] = ( -1 +1 -3 -1) = -4

We divide -4 by the number of sources to get the data sent by source B

-1 in CDMA represents data bit 0. So source B sent a 0.

To find the data sent by source C, we take product (Scalar product / dot product) of  the data on the channel with the chip sequence of source C

[-1 -1 -3 +1] . [+1 +1 -1 -1] = ( -1 -1 +3 -1) = 0

We divide 0 by the number of sources to get the data sent by source C

0 in CDMA represents silence, so source C did not send anything and was silent.